A heavy piece of hanging sculpture is suspended by a90-cm-long, 5.0 g steel wire. When the wind
blows hard, the wire hums at its fundamental frequency of 80Hz. What is the mass of the
sculpture?
Answer #1
Concepts and reason

The concepts required to solve this problem are linear density, tension in wire and fundamental frequency of the string.

To calculate the mass of the sculpture firstly calculate the tension in wire and then calculate the linear density.

Fundamentals

The general expression for fundamental frequency of the string is,

f=12LTμf = \frac{1}{{2L}}\sqrt {\frac{T}{\mu }}

Here, LLis length of the wire , μ\mu is the linear density and TTis the tension in the wire.

The expression for tension in wire is,

T=MgT = Mg

Here, MMis mass of the sculpture and ggis the acceleration due to gravity.

The expression for linear density is,

μ=mL\mu = \frac{m}{L}

Here, mmis mass of the wire.

The expression for linear density is,

μ=mL\mu = \frac{m}{L}

Substitute 5g5{\rm{ g}} for mmand 90cm90\,\,{\rm{cm}} forLL in above equation of linear density.

μ=5g(1kg103g)90cm(1m102cm)=5.555×103kgm1\begin{array}{c}\\\mu = \frac{{5{\rm{ g}}\left( {\frac{{1\;{\rm{kg}}}}{{{{10}^3}{\rm{ g}}}}} \right)}}{{{\rm{90 cm}}\left( {\frac{{1\;{\rm{m}}}}{{{{10}^2}\;{\rm{cm}}}}} \right)}}\\\\ = 5.555 \times {10^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}}\\\end{array}

The expression for tension in wire is,

T=MgT = Mg

Substitute 9.8ms29.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}} for ggin above equation tension.

T=M(9.8ms2)T = M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right)

The general expression for fundamental frequency of the string is,

f=12LTμf = \frac{1}{{2L}}\sqrt {\frac{T}{\mu }}

Substitute 5.555×103kgm15.555 \times {10^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}} for μ\mu , 90cm90\,\,{\rm{cm}} forLL and M(9.8ms2)M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right) for TTin above equation.

f=12(90cm(1m102cm))M(9.8ms2)5.555×103kgm1=0.555m1M(9.8ms2)5.555×103kgm1\begin{array}{c}\\f = \frac{1}{{2\left( {90\,\,{\rm{cm}}\left( {\frac{{1\;{\rm{m}}}}{{{{10}^2}\;{\rm{cm}}}}} \right)} \right)}}\sqrt {\frac{{M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right)}}{{5.555 \times {{10}^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}}}}} \\\\ = 0.555\;{{\rm{m}}^{ - 1}}\sqrt {\frac{{M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right)}}{{5.555 \times {{10}^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}}}}} \\\end{array}

Substitute 80Hz80{\rm{ Hz}} for ff in above equation.

80Hz0.555m1=M(9.8ms2)5.555×103kgm1(1764kg1m2s2)(144Hzm)2=M(9.8ms2)5.555×103kgm1M=11.76kg\begin{array}{c}\\\frac{{80{\rm{ Hz}}}}{{0.555\;{{\rm{m}}^{ - 1}}}} = \sqrt {\frac{{M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right)}}{{5.555 \times {{10}^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}}}}\left( {1764{\rm{ k}}{{\rm{g}}^{ - 1}} \cdot {{\rm{m}}^2} \cdot {{\rm{s}}^{ - 2}}} \right)} \\\\{\left( {144{\rm{ Hz}} \cdot {\rm{m}}} \right)^2} = \frac{{M\left( {9.8{\rm{ m}} \cdot {{\rm{s}}^{ - 2}}} \right)}}{{5.555 \times {{10}^{ - 3}}\;{\rm{kg}} \cdot {{\rm{m}}^{ - 1}}}}\\\\M = 11.76{\rm{ kg}}\\\end{array}

Ans:

The mass of the sculpture is 11.76kg11.76{\rm{ kg}}.